# Biased Minds Meet Biased AI: How Class Imbalance Shapes Appropriate Reliance and Interacts with Human Base Rate Neglect

### v1

@@ 4. Hypotheses
Based on previous results, RSR and RAIR were expected to differ between model conditions (Schemmer2023Appropriate). In particular, in the unbalanced condition, where much of the AI advice implies the same disease, users may find it increasingly difficult to discern whether an AI’s advice is correct or incorrect. This may be because repeated exposure to advice favouring the same outcome could reinforce the perception that the AI’s suggestion is invalid. Further, we hypothesized that RAIR would be more affected by class imbalance based on the pilot study. We assume that this effect is larger because RAIR requires users to change away from their initial decision. These change-decisions could be avoided disproportionally often, because they require more (cognitive) effort (chi2020inhibiting).
@@ 5.1. Dataset Creation
We created a synthetic dataset using symptom data for the common cold, influenza, and dengue from several reputable medical websites (i.e., MSDManualcdcMAYOCLINICMedlinePlus). Symptoms were standardized and merged where necessary to ensure consistency (e.g., ”scratchy or sore throat” was coded as “sore throat”), resulting in 33 unique symptoms across the three diseases. Symptom frequency ratings (0-4) were assigned based on how often they appeared across websites (0= never mentioned to 4 = mentioned on all websites), and these ratings were used to generate probabilities for each symptom’s presence. We generated 300,000 individual cases (100,000 per disease) and introduced noise to reflect realistic AI accuracy levels (above 80% (Vereschak2021How)). For details on the dataset and methodology, see the codebook on OSF. We selected ten symptoms for analysis based on variable importance and non-diagnosticity. Non-diagnosticity was crucial since ignoring base rates can sometimes be rational (e.g., uniquely associating haemorrhagic fever with dengue). However, Since 40%-80% of dengue cases are mild or asymptomatic, including non-diagnostic symptoms reflects real-world scenarios with ambiguous priors  (division_communicable_diseases_dengue_2024) (see Figure 3 for symptoms).
@@ 5.2. Model Training
The generated dataset was then used to train two machine learning models using Tidymodels (Kuhn2020Tidymodels:). After an initial split into training and testing sets (80% and 20%, respectively), the data was balanced to contain the same amount of cases for each disease and the same sample size as the following unbalanced dataset (n = 35,549, per disease). Subsequently, a first random forest model with 500 trees was specified using the ranger-package (Wright2017ranger:). The accuracy on each set was satisfactory to evaluate an AI-assisted decision-making tool (Vereschak2021How) - 82.8% for the training set and 81.9% for the testing set, constituting our balanced model. For the unbalanced model, the same testing set was chosen; however, to create class imbalance, only 5% of common cold and 5% of influenza cases were randomly drawn to remain in the training set, resulting in a total of 10’862 cases for common cold and influenza and 95’979 cases for dengue. Then, the same model specification as before was applied. This second model achieved an accuracy of 95.0% for the training set, which follows the theoretical assumption that accuracy is an inadequate measure when class imbalance is present (guo2008class; megahed2021class). The accuracy on the testing set was 73.2%, which shows that the model performance did, as expected, decrease for unbalanced data. To create the stimulus material, 30 correct and 30 incorrect predictions were randomly sampled from each model (120 cases across both models) following the procedure by Schemmer2023Appropriate. This method ensured that visible accuracy was held constant across conditions and allowed meaningful computation of appropriate reliance measures. Observed differences in reliance therefore reflect differences in the distribution of errors rather than global model quality. The unbalanced condition clearly over-sampled “dengue” cases, while the balanced condition maintained an even disease distribution (see Appendix).
@@ 5.4. Procedure
Figure 3. User interfaces presented to participants during the trials. Left is the pre-advice screen. Right is the AI advice screen.:
@@
Participants completed the study on the platform EFS Unipark by Tivian [^2] Participant data was fully anonymised, in line with the General Data Protection Regulation (GDPR). They first provided informed consent. After giving consent, they were given instructions for the task to be completed and presented with a cover story about a company (AVDTM) testing two different AI screening tools for disease detection in Europe. The design ensured participants were unaware of the class imbalance in the training data and the study’s purpose. This was intentional because it mirrors real-world conditions, where users rarely have access to model information. They were then provided information about the diseases, including verbal descriptions of their base rates. These verbal descriptions were constructed to render the base rates of the disease salient, making the descending order and size of their prevalence very clear. The common cold was described as ”most prevalent human disease” and that people typically have ”more than one infection a year”. Influenza was described as ”common” and that many individuals experience it ”more than once throughout their lives”. While dengue was described as ”very rare” and only ”six locally acquired cases were reported in Europe”. Then, participants were provided with a practice example to familiarize themselves with the task before the study began. Participants first decided on their own, followed by AI advice, and then had the opportunity to revise their initial decision. One trial consisted of a pre-advice decision and post-advice decision. Participants were initially presented with a set of symptoms [^3] Symptom order in the table was randomized at the survey’s start for each participant but remained fixed within the survey to avoid confusion. and asked to select which disease—common cold, influenza, or dengue—they believed the random individual was suffering from. After this initial decision, the tool showed the AI advice on diagnosis [^4] PPV estimates were also asked before and after advice.. They were then allowed to revise their initial decision. If they had forgotten the symptoms for the trial, they could access an info box to see them again. Participants completed 60 trials per model, totalling 120 trials, with decisions made both before and after receiving advice (see Figure 3 for an example trial). The order in which participants interacted with the models was determined randomly [^5] After finishing all trials, participants filled out questionnaires assessing numeracy (McNaughton2015Validation) and graph literacy (Okan2019Using), which were intended for a potential follow-up.. They then completed demographic questionnaires. Participants had the opportunity to provide feedback and received a debriefing on the true purpose of the study. Given that deceptions were employed, participants were also given another explicit chance to withdraw their consent. To ensure data quality, we followed recommendations by BRUHLMANN2020100022 and included two instructed response items (curran_methods_2016) and a question assessing self-reported data quality (meade_identifying_2012). A procedure flow-chart and the exact survey instructions can be accessed in the Appendix.
@@ 5.5. Measures
The dimensions of appropriateness of reliance are calculated using the formulae adopted from Schemmer2023Appropriate:
@@ 5.6. Participants
All main study participants were recruited from the crowd-sourcing platform Prolific, which has shown sufficient data quality in large scale studies (e.g., scharowski2024trustdistrusttrustmeasures; Perrig2024PXI; Perrig2023visAWI). In total, 48 participants aged 18 and over, residing in the United Kingdom (UK) participated. This population was chosen to ensure English proficiency. Participants were reimbursed £8.50, deemed ”fair pay” by Prolific guidelines (Prolific2023pay). Participants were informed of an accuracy bonus of up to £0.50, which all received regardless of performance. One participant was removed for requesting data exclusion, and another for interrupting the survey. After data cleaning, 46 participants remained in the sample. 23 identified as men, 23 as women. On average, participants were 37.5 years old ($SD=13.07,min=18,max=78$). 34 participants indicated tertiary, three vocational training, seven high school, one primary school and one preferred not to disclose their highest education. Participants reported an average medical expertise score 2.85 ($SD=1.55,min=1,max=7$) with modes at 1 and 2 (bimodal). AI expertise was slightly higher, with a mean of 3.22 ($SD=1.79,min=1,max=7$), and mode 1. Thus, the sample was, on average, composed of users. The sample size was primarily limited by resources, though power analysis confirmed sufficient sample size to detect a medium effect for all pre-registered analyses presented. Aiming for 80% power and an alpha of 0.05, analyses were conducted using G*Power Version 3.1.9.6 (faul_gpower_2007). Power analysis estimated that 34 participants would be sufficient to detect a medium effect (Cohen’s d = 0.50) in a two-tailed paired-sample t-test examining the impact of class imbalance on appropriate reliance. Although the analysis methods for some hypotheses later changed, the initial power analyses indicated a required sample size of 43 for the planned compound-bias analyses.
@@ 6. Results
Figure 4. Differences in appropriateness of reliance between model conditions. Grey lines show within-subject changes. Dashed lines divided the figure into four response quadrants at RAIR= 0.33 and RSR= 0.33 which indicate the chance level.:
@@ 6.1. Class Imbalance and the Appropriateness of Reliance
To test these descriptive differences, we conducted an asymptotic Wilcoxon signed‐rank test with continuity correction due to non‐normality (this test was pre-registered). RAIR increased significantly from unbalanced to balanced (Z = 5.93, p << .001), supporting H1a; RSR decreased significantly (Z=2.54,p=.011Z=-2.54,p=.011), supporting H1b. Effect sizes were large for RAIR (r = .85) and moderate for RSR (r=.40r=.40(cohen1992power), in line with H1c. Thus, all pre-registered hypotheses for the effects of class imbalance on appropriateness of reliance dimensions could be confirmed.
@@ 6.2. Compound Bias: Class Imbalance and Base Rate Neglect
Figure 5. Estimated frequencies of diseases under balanced and unbalanced conditions, before and after AI advice across trials. Error bars represent 95% Confidence intervals.:
@@
To test the observed patterns, we used multinomial logistic regression from the mclogit package (mclogit2022elff), setting the common cold as the reference category. We specified a series of increasingly complex fixed-effects models: intercept-only; main effects of advice and model balance; their interaction; addition of trial; and the three-way interaction (advice x model balance x trial). The full interaction model provided the best fit with the lowest Akaike information criterion (AIC) and significant likelihood-Ratio test (LRT) improvements. There was no evidence of overdispersion (residual deviance/df = 1.08) and acceptable multicollinearity (all variance inflation factors small, one moderate at VIF = 5.13)  (hairMultivariateDataAnalysis2014). We thus selected the full-interaction model with advice, model balance, and trial as interacting predictors (see Appendix for model selection details).
@@
Table 2. Overview of the hypotheses and their confirmation status. H1a-H1c were confirmatory, while H2a-H3b were exploratory due to amendments in analysis methods (Nosek2018prereg).
@@ 7.1. Effects of Class Imbalance on the Appropriateness of Reliance
The results demonstrated that users’ appropriateness of reliance changed depending on whether they interacted with biased or unbiased AI. When interacting with the balanced model, both RAIR and RSR were above 50% on average. This indicates that participants appropriately relied more often on the AI, changing their initially incorrect decisions and disregarding incorrect AI advice. In contrast, when interacting with the unbalanced model, RSR was higher than with the balanced model, indicating that participants more often relied on themselves and maintained their initial correct decisions. However, this came at the cost of having low RAIR (under 30%), showing that participants often disregarded the AI, even when its advice was correct (see Figure 4). Although counter-intuitive, this aligns with evidence that people can detect systematic biases (Schemmer2023Appropriate). If systematic biases, such as class imbalance are detected, some might stop relying on AI altogether (RSR=1 and RAIR=0). However, an AI affected by class imbalance can still be fairly accurate. Thus, some participants show under-reliance when they detect systematic errors. This reflects AI aversion, where observing algorithmic failure diminishes confidence in its predictions (dietvorst_algorithm_2015).
@@
Comparing the biased and unbiased AI, we observed that interacting with a biased AI had a more pronounced effect on RAIR than RSR. We expected this larger effect due to the higher costs associated with changing one’s decisions instead of sticking with the initial one (chi2020inhibiting). Crucially, users performed better than chance in discriminating between correct and incorrect AI advice when interacting with the balanced model. Past research has shown that people under-rely on AI even when explanations for its advice were provided (Schemmer2023Appropriate). This suggests that class imbalance may have a stronger influence on behaviour than AI explanations, supporting recent arguments that static AI explanations do not align well with human cognitive processes (miller2023xaidead; lai2020Why). The significant change between the two models, along with its effect size, makes this influence plausible, highlighting that class imbalance is not only a technical issue but also a critical bias that must be addressed to improve reliance calibration in AI-assisted decision-making. However, we do not advocate excluding explainability techniques for appropriate human-AI decision-making. Although the effects are smaller, RAIR appears to be influenced by explanations (Schemmer2023Appropriate). Moreover, diagnostic tools are also influenced by explanations (tsai2021symptomchecker). Building on this, a promising direction may be to combine measures that address model balance with improved AI explanations to further enhance human-AI decision-making. One promising explainability technique could be visualisation (Zehrung2021visexmachina; Yin2019accuracy; vanBerkel2021fairness). Tangible visualisations of training data could help users understand the sampling shortcomings of AI. Icon arrays, in particular, may be effective for displaying base rates, as they have been shown to reduce BRN and improve risk judgments in medical decision-making (Galesic2011Graph; witt_visual_2022).

### v2

@@ 4. Hypotheses
Based on previous results, RSR and RAIR were expected to differ between model conditions (Schemmer et al., 2023). In particular, in the unbalanced condition, where much of the AI advice implies the same disease, users may find it increasingly difficult to discern whether an AI’s advice is correct or incorrect. This may be because repeated exposure to advice favouring the same outcome could reinforce the perception that the AI’s suggestion is invalid. Further, we hypothesized that RAIR would be more affected by class imbalance based on the pilot study. We assume that this effect is larger because RAIR requires users to change away from their initial decision. These change-decisions could be avoided disproportionally often, because they require more (cognitive) effort (Chi et al., 2020).
@@ 5.1. Dataset Creation
We created a synthetic dataset using symptom data for the common cold, influenza, and dengue from several reputable medical websites (i.e., MSD Manual (2024)CDC (2024)Mayo Clinic (2024)MedlinePlus (2024)). Symptoms were standardized and merged where necessary to ensure consistency (e.g., ”scratchy or sore throat” was coded as “sore throat”), resulting in 33 unique symptoms across the three diseases. Symptom frequency ratings (0-4) were assigned based on how often they appeared across websites (0= never mentioned to 4 = mentioned on all websites), and these ratings were used to generate probabilities for each symptom’s presence. We generated 300,000 individual cases (100,000 per disease) and introduced noise to reflect realistic AI accuracy levels (above 80% (Vereschak et al., 2021)). For details on the dataset and methodology, see the codebook on OSF. We selected ten symptoms for analysis based on variable importance and non-diagnosticity. Non-diagnosticity was crucial since ignoring base rates can sometimes be rational (e.g., uniquely associating haemorrhagic fever with dengue). However, Since 40%-80% of dengue cases are mild or asymptomatic, including non-diagnostic symptoms reflects real-world scenarios with ambiguous priors  (Federal Office of Public Health, Division of Communicable Diseases, 2023) (see Figure 3 for symptoms).
@@ 5.2. Model Training
The generated dataset was then used to train two machine learning models using Tidymodels (Kuhn and Wickham, 2020). After an initial split into training and testing sets (80% and 20%, respectively), the data was balanced to contain the same amount of cases for each disease and the same sample size as the following unbalanced dataset (n = 35,549, per disease). Subsequently, a first random forest model with 500 trees was specified using the ranger-package (Wright and Ziegler, 2017). The accuracy on each set was satisfactory to evaluate an AI-assisted decision-making tool (Vereschak et al., 2021) - 82.8% for the training set and 81.9% for the testing set, constituting our balanced model. For the unbalanced model, the same testing set was chosen; however, to create class imbalance, only 5% of common cold and 5% of influenza cases were randomly drawn to remain in the training set, resulting in a total of 10’862 cases for common cold and influenza and 95’979 cases for dengue. Then, the same model specification as before was applied. This second model achieved an accuracy of 95.0% for the training set, which follows the theoretical assumption that accuracy is an inadequate measure when class imbalance is present (Guo et al., 2008; Megahed et al., 2021). The accuracy on the testing set was 73.2%, which shows that the model performance did, as expected, decrease for unbalanced data. To create the stimulus material, 30 correct and 30 incorrect predictions were randomly sampled from each model (120 cases across both models) following the procedure by Schemmer et al. (2023). This method ensured that visible accuracy was held constant across conditions and allowed meaningful computation of appropriate reliance measures. Observed differences in reliance therefore reflect differences in the distribution of errors rather than global model quality. The unbalanced condition clearly over-sampled “dengue” cases, while the balanced condition maintained an even disease distribution (see Appendix).
@@ 5.4. Procedure
Figure 3. User interfaces presented to participants during the trials. Left is the pre-advice screen. Right is the AI advice screen.
Two user interfaces. The left user interface shows a symptom list where the selected individual has symptoms such as runny nose, sneezing, and fatigue. The user is asked to choose between influenza, common cold, and dengue based on these symptoms and provide a rating on their estimated probability that the person suffers from this disease. The right image shows the user interface displaying the AI’s advice of ’common cold’ with a 98% probability. Below, the user is asked to choose a disease and adjust the their probability estimate based on the AI advice.
@@
Participants completed the study on the platform EFS Unipark by Tivian [^2] Participant data was fully anonymised, in line with the General Data Protection Regulation (GDPR). They first provided informed consent. After giving consent, they were given instructions for the task to be completed and presented with a cover story about a company (AVDTM) testing two different AI screening tools for disease detection in Europe. The design ensured participants were unaware of the class imbalance in the training data and the study’s purpose. This was intentional because it mirrors real-world conditions, where users rarely have access to model information. They were then provided information about the diseases, including verbal descriptions of their base rates. These verbal descriptions were constructed to render the base rates of the disease salient, making the descending order and size of their prevalence very clear. The common cold was described as ”most prevalent human disease” and that people typically have ”more than one infection a year”. Influenza was described as ”common” and that many individuals experience it ”more than once throughout their lives”. While dengue was described as ”very rare” and only ”six locally acquired cases were reported in Europe”. Then, participants were provided with a practice example to familiarize themselves with the task before the study began. Participants first decided on their own, followed by AI advice, and then had the opportunity to revise their initial decision. One trial consisted of a pre-advice decision and post-advice decision. Participants were initially presented with a set of symptoms [^3] Symptom order in the table was randomized at the survey’s start for each participant but remained fixed within the survey to avoid confusion. and asked to select which disease—common cold, influenza, or dengue—they believed the random individual was suffering from. After this initial decision, the tool showed the AI advice on diagnosis [^4] PPV estimates were also asked before and after advice.. They were then allowed to revise their initial decision. If they had forgotten the symptoms for the trial, they could access an info box to see them again. Participants completed 60 trials per model, totalling 120 trials, with decisions made both before and after receiving advice (see Figure 3 for an example trial). The order in which participants interacted with the models was determined randomly [^5] After finishing all trials, participants filled out questionnaires assessing numeracy (McNaughton et al., 2015) and graph literacy (Okan et al., 2019), which were intended for a potential follow-up.. They then completed demographic questionnaires. Participants had the opportunity to provide feedback and received a debriefing on the true purpose of the study. Given that deceptions were employed, participants were also given another explicit chance to withdraw their consent. To ensure data quality, we followed recommendations by Brühlmann et al. (2020) and included two instructed response items (Curran, 2016) and a question assessing self-reported data quality (Meade and Craig, 2012). A procedure flow-chart and the exact survey instructions can be accessed in the Appendix.
@@ 5.5. Measures
The dimensions of appropriateness of reliance are calculated using the formulae adopted from Schemmer et al. (2023):
@@ 5.6. Participants
All main study participants were recruited from the crowd-sourcing platform Prolific, which has shown sufficient data quality in large scale studies (e.g., Scharowski et al. (2024); Perrig et al. (2024a, b)). In total, 48 participants aged 18 and over, residing in the United Kingdom (UK) participated. This population was chosen to ensure English proficiency. Participants were reimbursed £8.50, deemed ”fair pay” by Prolific guidelines (Prolific, 2023). Participants were informed of an accuracy bonus of up to £0.50, which all received regardless of performance. One participant was removed for requesting data exclusion, and another for interrupting the survey. After data cleaning, 46 participants remained in the sample. 23 identified as men, 23 as women. On average, participants were 37.5 years old ($SD=13.07,min=18,max=78$). 34 participants indicated tertiary, three vocational training, seven high school, one primary school and one preferred not to disclose their highest education. Participants reported an average medical expertise score 2.85 ($SD=1.55,min=1,max=7$) with modes at 1 and 2 (bimodal). AI expertise was slightly higher, with a mean of 3.22 ($SD=1.79,min=1,max=7$), and mode 1. Thus, the sample was, on average, composed of users. The sample size was primarily limited by resources, though power analysis confirmed sufficient sample size to detect a medium effect for all pre-registered analyses presented. Aiming for 80% power and an alpha of 0.05, analyses were conducted using G*Power Version 3.1.9.6 (Faul et al., 2007). Power analysis estimated that 34 participants would be sufficient to detect a medium effect (Cohen’s d = 0.50) in a two-tailed paired-sample t-test examining the impact of class imbalance on appropriate reliance. Although the analysis methods for some hypotheses later changed, the initial power analyses indicated a required sample size of 43 for the planned compound-bias analyses.
@@ 6. Results
Figure 4. Differences in appropriateness of reliance between model conditions. Grey lines show within-subject changes. Dashed lines divided the figure into four response quadrants at RAIR= 0.33 and RSR= 0.33 which indicate the chance level.
Scatter plot showing the relationship between Relative AI-Reliance (RAIR) on the x-axis and Relative Self-Reliance (RSR) on the y-axis, with values ranging from 0 to 1. Mean value for the balanced condition is RSR: 0.58 and RAIR: 0.55. Whereas the unbalanced condition means are RSR: 0.75 and RAIR: 0.12. Showing that users self-relied more in the unbalanced condition and relied less on AI. In the balanced condition this changed to people on average relying more appropriately.
@@ 6.1. Class Imbalance and the Appropriateness of Reliance
To test these descriptive differences, we conducted an asymptotic Wilcoxon signed‐rank test with continuity correction due to non‐normality (this test was pre-registered). RAIR increased significantly from unbalanced to balanced (Z = 5.93, p << .001), supporting H1a; RSR decreased significantly (Z=2.54,p=.011Z=-2.54,p=.011), supporting H1b. Effect sizes were large for RAIR (r = .85) and moderate for RSR (r=.40r=.40(Cohen, 1992), in line with H1c. Thus, all pre-registered hypotheses for the effects of class imbalance on appropriateness of reliance dimensions could be confirmed.
@@ 6.2. Compound Bias: Class Imbalance and Base Rate Neglect
Figure 5. Estimated frequencies of diseases under balanced and unbalanced conditions, before and after AI advice across trials. Error bars represent 95% Confidence intervals.
Four line plots illustrate the trajectories of participants’ mean estimated frequencies for three diseases—common cold, influenza, and dengue—under balanced and unbalanced conditions, before and after receiving advice. In the top-left plot (pre-advice, balanced), dengue estimates are consistently low, while estimates for common cold increase across trials, and influenza shows an interaction with common cold and decreases. In the bottom-left plot (pre-advice, unbalanced), dengue estimates increase, while common cold and influenza decrease. In the top-right plot (post-advice, balanced), dengue estimates remain low, while common cold estimates rise and diverge further from influenza. In the bottom-right plot (post-advice, unbalanced), dengue estimates were higher to begin with and slightly increases across trials, while common cold and influenza estimates converge and decline slightly.
@@
To test the observed patterns, we used multinomial logistic regression from the mclogit package (Elff, 2022), setting the common cold as the reference category. We specified a series of increasingly complex fixed-effects models: intercept-only; main effects of advice and model balance; their interaction; addition of trial; and the three-way interaction (advice x model balance x trial). The full interaction model provided the best fit with the lowest Akaike information criterion (AIC) and significant likelihood-Ratio test (LRT) improvements. There was no evidence of overdispersion (residual deviance/df = 1.08) and acceptable multicollinearity (all variance inflation factors small, one moderate at VIF = 5.13)  (Hair et al., 2014). We thus selected the full-interaction model with advice, model balance, and trial as interacting predictors (see Appendix for model selection details).
@@
Table 2. Overview of the hypotheses and their confirmation status. H1a-H1c were confirmatory, while H2a-H3b were exploratory due to amendments in analysis methods (Nosek et al., 2018).
@@ 7.1. Effects of Class Imbalance on the Appropriateness of Reliance
The results demonstrated that users’ appropriateness of reliance changed depending on whether they interacted with biased or unbiased AI. When interacting with the balanced model, both RAIR and RSR were above 50% on average. This indicates that participants appropriately relied more often on the AI, changing their initially incorrect decisions and disregarding incorrect AI advice. In contrast, when interacting with the unbalanced model, RSR was higher than with the balanced model, indicating that participants more often relied on themselves and maintained their initial correct decisions. However, this came at the cost of having low RAIR (under 30%), showing that participants often disregarded the AI, even when its advice was correct (see Figure 4). Although counter-intuitive, this aligns with evidence that people can detect systematic biases (Schemmer et al., 2023). If systematic biases, such as class imbalance are detected, some might stop relying on AI altogether (RSR=1 and RAIR=0). However, an AI affected by class imbalance can still be fairly accurate. Thus, some participants show under-reliance when they detect systematic errors. This reflects AI aversion, where observing algorithmic failure diminishes confidence in its predictions (Dietvorst et al., 2015).
@@
Comparing the biased and unbiased AI, we observed that interacting with a biased AI had a more pronounced effect on RAIR than RSR. We expected this larger effect due to the higher costs associated with changing one’s decisions instead of sticking with the initial one (Chi et al., 2020). Crucially, users performed better than chance in discriminating between correct and incorrect AI advice when interacting with the balanced model. Past research has shown that people under-rely on AI even when explanations for its advice were provided (Schemmer et al., 2023). This suggests that class imbalance may have a stronger influence on behaviour than AI explanations, supporting recent arguments that static AI explanations do not align well with human cognitive processes (Miller, 2023; Lai et al., 2020). The significant change between the two models, along with its effect size, makes this influence plausible, highlighting that class imbalance is not only a technical issue but also a critical bias that must be addressed to improve reliance calibration in AI-assisted decision-making. However, we do not advocate excluding explainability techniques for appropriate human-AI decision-making. Although the effects are smaller, RAIR appears to be influenced by explanations (Schemmer et al., 2023). Moreover, diagnostic tools are also influenced by explanations (Tsai et al., 2021). Building on this, a promising direction may be to combine measures that address model balance with improved AI explanations to further enhance human-AI decision-making. One promising explainability technique could be visualisation (Zehrung et al., 2021; Yin et al., 2019; van Berkel et al., 2021). Tangible visualisations of training data could help users understand the sampling shortcomings of AI. Icon arrays, in particular, may be effective for displaying base rates, as they have been shown to reduce BRN and improve risk judgments in medical decision-making (Galesic and Garcia-Retamero, 2011; Witt and Dhami, 2022).